You are a participant in a demolition derby. The d
Answer #1

Kinetic Energy = 1/2 * mv^2

(a)
If we Increase the speed of Car 5 times.

The kinetic Energy = 1/2 * m(5v)^2
K.E = 1/2 * mv^2 * 25
K.Enew = 25 * K.Eold
So Kinetic Energy will increase 25 fold.


(b)

If we do double the speed,
K.E = 1/2 * m (2v)^2
K.E = 4 * 1/2 * mv^2

i.e Kinetic Energy will be 4 times the old K.E.
If we triple the mass, KInetic Energy will be 3 times the old K.E.

So, Doubling the Speed will be better.

(c)
Kinetic Energy will be lost into some other form of energy which can be Thermal Energy, Sound , Or the other object will gain some Kinetic Energy due to the crash.